OpenStudy (anonymous):

Please Help The square of a number decreased by 25 is -9 A. Find the x-intercepts. B. Find the vertex of the graph. C. What shape is the graph? Describe what you see.

OpenStudy (anonymous):

I just helped someone with this a couple days ago, so this should be easy :)

OpenStudy (anonymous):

The first one is asking you- Something Squared minus 25 equals -9

OpenStudy (anonymous):

Your equation \[n^2-25=-9\]

OpenStudy (anonymous):

Now you have to solve for n

OpenStudy (anonymous):

I'm having problems with, part B and C

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm n^2-25=-9 }\) \(\large\color{blue}{ \rm n^2-25+9=-9 +9}\) \(\large\color{blue}{ \rm n^2-16=0}\) \(\large\color{blue}{ \rm (n-4)(n+4)=0}\) \(\large\color{blue}{ \rm n=~±~4}\)

OpenStudy (anonymous):

How would you graph that?

OpenStudy (solomonzelman):

graph the solutions? Graph (4,0) and (-4,0)